# Explanation Bias is a Product: Revealing the Hidden Lexical and Position Preferences in Post-Hoc Feature Attribution

### v1

@@ 3.5 Results
As expected, our fine-tuned models achieve chance-level f1-scores, in the range .33–.55. We report the results for all bias metrics based on the top-1 attributions across random seeds for each model type and attribution method. We present the full visuals of the bias distributions in Appendix A.4.
@@ B ModB B ModB B ModB B ModB B ModB B ModB
Position Bias
PartSHAP .19 .22 .26 .26 .23 .10 .13 .15 .21 .26 .13 .09*
@@
Lexical Bias
PartSHAP .09 .03 .02 .05 .19 .22 .11 .20 .02 .05 .19 .30
@@
We compare Bias-attr results to Bias-agg results to determine whether the difference between methods is indicative of a difference between the methods and a no-bias-baseline. For this, we consider the highest Bias-attr method per dataset–model pair: in 6/6 cases it corresponds to the highest Bias-agg method for position and in 4/6 cases for lexical bias. If we look at low Bias-attr scores, this parallel pattern with Bias-agg pattern is less clear: several methods exhibit comparably low scores. This indicates that, at least in our experiments, methods with a stronger bias compared to one another (Bias-attr) are also more biased overall compared to a uniform baseline (Bias-agg).
@@ 4 Experiments on Causal Relation Data
Causal relations can be expressed in many ways, both explicitly and implicitly solstadbott. Here we use data with explicit markers, such as causal verbs (cause, depend on) and adverbial phrases (because, as a result of). These strong markers make this task particularly useful for our purpose, contrasting with the artificial dataset where there are no overt markers the models can learn from. There are two main tasks in the domain of causality detection: sentence classification (does it contain causal information?) and causal span detection (cause, marker, and effect) (asghar2016, e.g.). We address the former task.
@@ 4.1 Data
The data is adapted from veldhuisetal, and consists of annotated texts that contain causal relations used for system dynamics modelling. They stem from documents on climate change, putinism policy and traffic congestion (sterman2003climatechange; mcfaul2020putin; sterman2000trafficcongestion). We finetune the classifiers on alterations of these texts to i) emulate the repetition of causal signals in the same input and ii) to control for position bias compared to a baseline. For this, we first randomly sample sentence triples (900 for training, 100 for testing). We then construct input instances as permutations of these triples, with each triple resulting in 6 different orderings of the three different sentences (5400 for training, 600 for testing). Positive instances (50% of the data) consist of concatenations of three different positive sentences (+++); the same logic applies to negative instances (---). The three sentences in +++ represent causal occurrences at beginning-, middle- and end-of-text. We name this diagnostic dataset causal. The task is to predict a positive or negative class label for each instance. We train and test on the perturbations so that the model should not learn that the dominant causal signals (each instance is expected to contain one) are tied to specific sentence positions. By only changing the order of the sentences (and not e.g. the syntactic structure) and keeping the tokens constant per permutation batch, we control for confounding factors that would arise from feeding different inputs. On average, positive sentences contain 24.6 words (6–60), negative 18.4 (1–57), while +++ and --- test instances contain 64.1 (22–128). Example inputs are given in Appendix A.1.
@@ 4.2 Bias and Baseline Distributions
To deepen our analysis, we differentiate between positive class attributions and negative class attributions. The former are expected to target strong causal signals, whereas the latter must deal with a signal absence, which makes attribution intuitively more challenging (e.g. dela2024none).
@@ 4.3 Results
BERT and ModernBERT fine-tuned on the causal task achieve f1-scores $\scriptstyle>$.98, hence forming an interesting counter scenario to the artificial experiments (f1-score $\scriptstyle\sim$.50). From a first lexical insight, we notice that targeted top-1 tokens vary across models and methods. For example, BERT’s LIME selects mostly causal verbs such as produce, increases, achieve, influence, depends, weaken, triggered. ModernBERT’s IntGrad selects a mix of different lexical elements, such as causal verbs (made, change, developed), but also punctuation, determiners and topic-related words (temperature, climate, regime, emissions). The latter implausible explanations increase the relevance of investigating where in the input they occur. Moreover, methods attribute more function words and punctuation in negative examples, making up for the absence of causal signals. In this section, we report the sentence position biases from these models.
@@
The scores in Figure 3 are generally low (.03–.15 across models): attribution methods yield a relatively consistent bias across sentence positions. This means that when a task is learnt successfully, a dominant causal signal is most likely identified irrespective of position. Low scores are in part due to a relatively low number of sentence positions (3), facilitating the task by conflating token attributions into sentence attributions. We notice that whereas LIME and IntGrad were roughly the most consistent on the artificial data, only LIME is on causal, suggesting that LIME is only lightly affected by the models’ different random seeds. Instead, the least consistent methods on the artificial sets (VanGrad for BERT, IntGrad × I for ModernBERT) are not the least on causal.
Different from the artificial setting, ModernBERT is not predominantly less biased than BERT, despite BERT’s high bias on IntGrad and IntGrad × I (.13 and.15). Finally, positive class biases are more consistent for BERT, while negative examples are marginally more consistent for ModernBERT.
@@
For BERT, the attributions point to the dominant (non-)causal signal appearing with a slight bias in the last sentence of each +++ or --- instance (Figure 4). This suggests that BERT is most biased towards late information in the input, especially on negative examples, which reflects the results on artificial data (Figures 5, 6, 7, Appendix A).
@@ 5 Conclusion and Future Directions
We proposed a bias evaluation framework on both artificial and natural language data, finding systematic preferences for lexical and positional information. We questioned how lexical and position bias relate and whether a recent BERT transformer would be less biased than its traditional version. We conclude that the two types of biases do not imply one another and that ModernBERT is not necessarily less biased than BERT. We also questioned whether evidence from anomalous explainers is more likely to be incorrect. To this end, we found repeated empirical evidence that methods differing most from other methods are also the most biased. Since our settings controlled for lexical and position confounders, this bias is expected to stem from the method itself and not from the backbone model.
@@
Acknowledgments
This research is funded by the Dutch National Science Organisation (NWO) through project code NWA.1292.19.399 – InDeep: Interpreting Deep Learning Models for Text and Sound. We would also like to thank TNO and the APPL.AI programme for their support of this project.
@@ A.2 Hardware and Runtimes
Table 3: Model fine-tuning times (5 epochs) on artificial and causal data along with explanation generation times per instance (6 default methods in Ferret (attanasio-etal-2023-ferret) with default parameters).
@@ A.3 Data Artifacts
We use the existing dataset by veldhuisetal. It was obtained on request to the authors, who agreed to publish the data splits used in our experiments on our repository. We advice to contact the authors for re-use of the data, intended for research purposes. The data was manually checked for offensive content or personal information, which were not encountered.
A.4 Complementary Visualisations

### v2

@@ 3.5 Results
As expected for this fully controlled experimental setting, our fine-tuned models achieve chance-level f1-scores, in the range .33–.55.333Importantly, the resulting biases are non-random, as we extensively discuss in Appendix A.4. In the next paragraphs, we report the results for all bias metrics based on the top-1 attributions across random seeds for each model type and attribution method. We plot the full bias distributions in Appendix A.5.
@@ B ModB B ModB B ModB B ModB B ModB B ModB
Position Bias PartSHAP .19 .22 .26 .26 .23 .10 .13 .15 .21 .26 .13 .09*
@@
Lexical Bias PartSHAP .09 .03 .02 .05 .19 .22 .11 .20 .02 .05 .19 .30
@@
We compare Bias-attr results to Bias-agg results to determine whether the difference between methods is indicative of a difference between the methods and a no-bias-baseline. For this, we consider the highest Bias-attr method per dataset–model pair: in 6/6 cases it corresponds to the highest Bias-agg method for position and in 4/6 cases for lexical bias. If we look at low Bias-attr scores, this parallel pattern with Bias-agg is less clear: several methods exhibit comparably low scores. This indicates that, at least in our experiments, methods with a stronger bias compared to one another (Bias-attr) are also more biased overall compared to a uniform baseline (Bias-agg).
@@ 4 Experiments on Causal Relation Data
Causal relations can be expressed in many ways, both explicitly and implicitly Solstad and Bott (2017). Here we use data with explicit markers, such as causal verbs (cause, depend on) and adverbial phrases (because, as a result of). These strong markers make this task particularly useful for our purpose, contrasting with the artificial dataset where there are no overt markers the models can learn from. There are two main tasks in the domain of causality detection: sentence classification (does it contain causal information?) and causal span detection (cause, marker, and effect) (Asghar, 2016, e.g.). We address the former task.
@@ 4.1 Data
The data is adapted from Veldhuis et al. (2024), and consists of annotated texts on causal relations used for system dynamics modelling. Topics include climate change, putinism policy and traffic congestion (McFaul, 2020, e.g.). We finetune the classifiers on alterations of these texts to i) emulate the repetition of causal signals in the same input and ii) to control for position bias compared to a baseline. For this, we first randomly sample sentence triples (900 for training, 100 for testing). We then construct input instances as permutations of these triples, with each triple resulting in 6 different orderings of the three different sentences (5400 for training, 600 for testing). Positive instances (50% of the data) consist of concatenations of three different positive sentences (+++); the same logic applies to negative instances (---). The three sentences in +++ represent causal occurrences at beginning-, middle- and end-of-text. We name this diagnostic dataset causal. The task is to predict a positive or negative class label for each instance. We train and test on the perturbations so that the model should not learn that the dominant causal signals (each instance is expected to contain one) are tied to specific sentence positions. By only changing the order of the sentences (and not e.g. the syntactic structure) and keeping the tokens constant per permutation batch, we control for confounding factors that would arise from feeding different inputs. On average, positive sentences contain 24.6 words (6–60), negative 18.4 (1–57), while +++ and --- test instances contain 64.1 (22–128). We provide examples in A.1.
@@ 4.2 Bias and Baseline Distributions
To deepen our analysis, we differentiate between positive class attributions and negative class attributions. The former are expected to target strong causal signals, whereas the latter must deal with a signal absence, which makes attribution intuitively more challenging (e.g. dela Cruz et al., 2024).
@@ 4.3 Results
BERT and ModernBERT fine-tuned on the causal task achieve f1-scores $\scriptstyle>$.98, hence forming an interesting counter scenario to the artificial experiments (f1-score $\scriptstyle\sim$.50). As a first lexical insight, the targeted top-1 tokens vary across models and methods. For example, BERT’s LIME selects mostly causal verbs such as produce, increases, achieve, influence, weaken. ModernBERT’s IntGrad selects a mix of different lexical elements, such as causal verbs (made, change, developed), but also punctuation, determiners and topic-related words (temperature, climate, regime, emissions). The latter implausible explanations increase the relevance of investigating where in the input they occur. Moreover, methods attribute more function words and punctuation in negative examples, making up for the absence of causal signals. In this section, we report the sentence position biases from these models.
@@
Figure 3 shows that attribution methods yield a relatively consistent bias across sentence positions (.03–.15 across models). This suggests that a dominant causal signal tends to be identified consistently across seeds irrespective of position. Low scores are in part due to a relatively low number of sentence positions (3), facilitating the task by conflating token attributions into sentence attributions. While LIME and IntGrad were the most consistent on the artificial data, only LIME is also the most consistent on causal, suggesting that LIME is only lightly affected by the seed. Different from the artificial setting, ModernBERT is not predominantly less biased than BERT, despite BERT’s higher bias on IntGrad and IntGrad × I (.13 and.15).
@@
For BERT, the attributions point to the dominant (non-)causal signal appearing with a slight bias in the last sentence of each +++ or --- instance (Figure 4). This suggests that BERT is most biased towards late information in the input, especially on negative examples, which is in line with the results on artificial data (Figures 5, 6, 7, Appendix A).
@@ 5 Conclusion and Future Directions
We proposed a bias evaluation framework on both artificial and natural language data, finding systematic preferences for lexical and positional information. We investigated how lexical and position bias relate and whether a recent BERT transformer would be less biased than its traditional version. We conclude that there is a trade-off between the two types of bias and that ModernBERT is not necessarily less biased than BERT. We also investigated whether evidence from anomalous explainers is more likely to be incorrect. To this end, we found repeated empirical evidence that methods differing most from other methods are also the most biased. Since our settings controlled for lexical and position confounders, this bias is expected to stem from the method itself and not from the backbone model.
@@ A.2 Hardware and Runtimes
Table 3: Model fine-tuning times (5 epochs) on artificial and causal data along with explanation generation times per instance (6 default methods in Ferret (Attanasio et al., 2023) with default parameters).
@@ A.3 Data Artifacts
We use the existing dataset by Veldhuis et al. (2024). It was obtained on request to the authors, who agreed to publish the data splits used in our experiments on our repository. We advice to contact the authors for re-use of the data, intended for research purposes. The data was manually checked for offensive content or personal information, which were not encountered.
A.4 Randomised Settings but Non-random Explanations
Explanation evaluation can be problematic when the test data is out-of-distribution (OOD) with respect to the training data (Hase et al., 2021). This does not apply to our setting, as we sample the training and test instances from the same data distribution. While the artificial instances might be unlikely or rare in pre-training, the highly frequent tokens that compose them (e.g. ‘.’, ‘,’) have regular pre-trained word embeddings and should therefore not be OOD with respect to the pre-training data. Technically, the latter OOD condition could even benefit our setup: as the artificial experiments target feature attribution bias irrespective of the data, the modelling of OOD examples would make it less likely to be influenced by pre-training data cues. Still, to rule out possible random noise in the top-1 frequency distributions caused by (improbable) OOD training effects, we report the results of a statistical randomisation test. This test should support the repeated nature of our fully controlled experimental setting, where explanations on thousands of examples should cancel out any random noise in the top-1 explanation distributions while highlighting any bias consistencies.
[…]
A.5 Complementary Visualisations